a car accelerates with a constant acceleration of 1.2ms^-1 along a long straight road.how long does it take the car to accelerate from 100kph to 110kph,and how far does it travel while doing so?
1
Expert's answer
2020-04-10T13:41:05-0400

Given:-

Acceleration(a) =1.2 m/s2

Car speed(accelerate from).

"u=100 kph = 72.78 m/s.\\[5pt]

v = 110 kph = 30.56 m/s."


Now,

From Kinetic equation.

"v=u+at\\[5pt]
\\t=\frac{v-u}{a}\\[8pt]
\\t=\frac{30..56 m/s-27.78 m/s}{1.2 m/s^{2}}\\[5pt]
t=2.315 sec."

So,

The time taken by car to travel from speed .

"\\t=2.315 sec."


Now,

Again , use following kinetic equation .

"v^{2}-u^{2}=2as\\[5pt]
s=\frac{v^{2}-u^{2}}{2a}\\[5pt]
s=\frac{(30.56m/s)^{2}-(27.78m/s)^{2}}{2\times 1.2m/s^{2}}\\[5pt]
s=67.56 m"

So,

The distance traveled by the the car is.

"s=67.56 m"